# Network Alignment with Transferable Graph Autoencoders

### v3

@@ Abstract
Network alignment is the task of establishing one-to-one correspondences between the nodes of different graphs and finds a plethora of applications in high-impact domains. However, this task is known to be NP-hard in its general form, and existing algorithms do not scale up as the size of the graphs increases. To tackle both challenges we propose a novel generalized graph autoencoder architecture, designed to extract powerful and robust node embeddings, that are tailored to the alignment task. We prove that the generated embeddings are associated with the eigenvalues and eigenvectors of the graphs and can achieve more accurate alignment compared to classical spectral methods. Our proposed framework also leverages transfer learning and data augmentation to achieve efficient network alignment at a large scale without retraining. Extensive experiments on both network and sub-network alignment with real-world graphs provide corroborating evidence supporting the effectiveness and scalability of the proposed approach.
@@ 1 Introduction
Network alignment, also known as graph matching, is a classical problem in graph theory, that aims to find node correspondence across different graphs and is vital in a number of high-impact domains (Emmert-Streib et al., 2016). In social networks, for instance, network alignment has been used for user deanonymization (Nilizadeh et al., 2014) and analysis (Ogaard et al., 2013), while in bioinformatics it is a key tool to identify functionalities in protein complexes (Singh et al., 2008), or to identify gene–drug modules (Chen et al., 2018). Graph matching also finds application in computer vision (Conte et al., 2003), sociology (Racz and Sridhar, 2021), or politics (Li et al., 2022), to name a few. Graph matching can be cast as a quadratic assignment problem (QAP), which is in general NP-hard
@@
Various approaches have been developed to tackle network alignment and can be divided into two main categories; i) optimization algorithms that attempt to approximate the QAP problem by relaxing the combinatorial constraints, ii) embedding methods that approach the problem by implicitly or explicitly generating powerful node embeddings that facilitate the alignment task. Optimization approaches, as (Anstreicher and Brixius, 2001; Vogelstein et al., 2015) employ quadratic programming relaxations, while (Klau, 2009) and (Peng et al., 2010) utilize semidefinite or Lagrangian-based relaxations respectively. Successive convex approximations were also proposed by (Konar and Sidiropoulos, 2020) to handle the QAP. Challenges associated with these methods include high computational cost, infeasible solutions, or nearly optimal initialization requirements. Embedding methods, on the other hand, overcome these challenges, but they usually produce inferior solutions, due to an inherent trade-off between embedding permutation-equivariance and the ability to capture the structural information of the graph. Typical embedding techniques include spectral and factorization methods (Umeyama, 1988; Feizi et al., 2019; Zhang and Tong, 2016; Kanatsoulis and Sidiropoulos, 2022), structural feature engineering methods (Berlingerio et al., 2013; Heimann et al., 2018), and random walk approaches (Perozzi et al., 2014; Grover and Leskovec, 2016a). Recently (Chen et al., 2020; Karakasis et al., 2021) have proposed joint node embedding and network alignment, to overcome these challenges, but these methods do not scale up as the size of the graph increases.
Figure 1: The overall architecture of a three-layer T-GAE and its training paradigm: The input signal is processed by a local MLP and then processed by GNN layers with skip connections. The outputs of all neural network layers are encoded by an MLP followed by a single-layer decoder to generate the reconstructed graph. The whole model is trained end-to-end with a single or multiple graphs.
Graph Neural Networks (GNNs) are powerful architectures that learn graph representations (embeddings). They have shown state-of-the-art performance in several tasks, including biology (Gainza et al., 2020; Strokach et al., 2020; Jiang et al., 2021), quantum chemistry (Gilmer et al., 2017), social networks and recommender systems (Ying et al., 2018; Wu et al., 2020; Liu et al., 2023; Yang et al., 2024). Recently, (Gao et al., 2021) proposed a GNN approach to match attributed graphs. The method used a joint embedding framework for pairs of graphs and achieved high levels of matching accuracy. However, this method does not scale to large graphs, since training graphs with large sizes is computationally prohibitive.
To address these challenges, we propose a novel self-supervised GNN framework to perform network alignment on a large scale. Specifically, we design a generalized transferable graph autoencoder (T-GAE) (shown in Fig. 1), to produce permutation equivariant and highly expressive embeddings, overcoming the challenges of other embedding techniques. T-GAE is trained on multiple graphs and learns node representations which are tailored to perform alignment between nodes of different graphs. The T-GAE representations combine the eigenvectors of the graph in a nonlinear fashion and are provably at least as good in network alignment as certain spectral methods. Additionally, the proposed framework leverages transfer learning and data augmentation to efficiently operate with large graphs. Training is performed with small graphs, in a self-supervised manner, and the trained encoder can be executed on large graphs. Extensive experiments with real-world benchmarks test the effectiveness and limits of the proposed T-GAE approach in the tasks of graph and sub-graph matching. The experimental results provide corroborating evidence that T-GAE offers an elegant framework for large-scale network alignment. Our contributions are summarized as follows:
  • (C1)
    We propose T-GAE, a generalized graph autoencoder architecture that can be trained with multiple graphs and produce expressive/permutation equivariant representations, tailored to network alignment.
  • (C2)
    We draw the the connection between T-GAE and spectral methods and prove that T-GAE is at least as good in graph matching as the absolute value of the graph eigenvectors.
  • (C3)
    We leverage data augmentation and transfer learning, to develop a robust encoder model and training framework that efficiently performs large scale network alignment.
  • (C4)
    We demonstrate the effectiveness and scalability of the proposed T-GAE with real-world, benchmark graphs in challenging graph and sub-graph matching settings.
@@ 2 Preliminaries
Graphs are represented by $\mathcal{G}:=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{1,\dots,N\}$ is the set of vertices (nodes) and =$\mathcal{E}=\left\{\left(v,u\right)\right\}$ correspond to edges between pairs of vertices. A graph is represented in a matrix form by a graph operator $\bm{S}\in\mathbb{R}^{N\times N}$, where $\bm{S}(i,j)$ quantifies the relation between node $i$ and node $j$ and $N=|\mathcal{V}|$ is the total number of vertices. In this work, we use the graph adjacency and normalized graph adjacency. Oftentimes, the nodes of the graph are associated with graph signals or node attributes $\bm{X}\in\mathbb{R}^{N\times D}$, that encode additional information about the nodes. In this paper, we study both network alignment of graphs with or without attributes.
@@ 2.1 Network Alignment
Figure 2: Robust GAE training algorithm as illustrated in Equation 8. For each of the training graph, we generate a number of augmented samples by randomly adding or removing edges. The node embedding of these augmented samples are decoded non-parametrically, and compared with the corresponding original graph to train the T-GAE encoder.
@@ 2.2 Spectral Decomposition of the Graph
where $\bm{E}\in\mathbb{R}^{N\times F}$ is embedding matrix and $\bm{E}[i,:]$ is the vector representation of node $i$. The optimization problem in (2) is a linear assignment problem and can be optimally solved in $\mathcal{O}\left(N^{3}\right)$ by the Hungarian method (Kuhn, 1955b). Simpler sub-optimal alternatives also exist that operate with $\mathcal{O}\left(N^{2}\right)$ or $\mathcal{O}\left(N~{}\text{log}(N)\right)$ flops.
@@ 3 Graph Neural Networks (GNNs) Upper-Bounds Spectral Methods for Network Alignment
The above property is not satisfied by other spectral methods. GNNs are also stable (Gama et al., 2020; Parada-Mayorga et al., 2023; Ruiz et al., 2021), transferable (Ruiz et al., 2020), and have high expressive power (Xu et al., 2019b; Abboud et al., 2021; Kanatsoulis and Ribeiro, 2022).
@@ 3.1 GNNs and Network Alignment
To characterize the ability of a GNN to perform network alignment we first point out the GNNs perform nonlinear spectral operations. Details can be found in Appendix B. We can prove that:
@@
The proof can be found in Appendix C. The assumption that the graph adjacencies have different eigenvalues is not restrictive. Real nonisomorphic graphs have different eigenvalues with very high probability (Haemers and Spence, 2004). Theorem 3.2 compares the network alignment power of a GNN with that of a spectral algorithm Umeyama (1988), that uses the absolute values of graph adjacency eigenvectors to match two different graphs. According to Theorem 3.2 there always exists a GNN that can perform at least as well as the spectral approach. The proof studies a GNN with white random input and measures the variance of the filter output. Then it shows that GNN layers are able to compute the absolute values of the graph adjacency eigenvectors when the adjacency has non-repeated eigenvalues. As a result there always exists a single layer GNN that outputs the same node features as the ones used in Umeyama (1988), which concludes our proof.
@@ 4 Proposed Method
We now leverage the favorable properties of GNNs (permutation equivariance, expressivity, and transferability) and design a GNN approach to tackle network alignment at a large-scale. Our approach learns low-dimensional node embeddings (Eq. 4) that enable graph matching via solving the linear assignment in (2) rather than a quadratic assignment problem in (1). In this section, we design a robust GNN framework such that the node embeddings are expressive to accurately match similar nodes and also stable to graph perturbations, so that they yield high-quality network alignment.
4.1 Learning geometry preserving embeddings
A fundamental property of node embeddings is to preserve the geometry and topological characteristics of the network. This will allow expressive node representations that can effectively approximate the original problem in (1) with the problem in (2). To achieve this goal we leverage an auto-encoder architecture that reconstructs the original graph from the node embeddings. Results on GNN expressivity indicate that this reconstruction is doable under specific conditions (Abboud et al., 2021). To build topology-preserving embeddings we solve the following optimization problem:
$\min_{\mathcal{H}}~{}l\left(~{}\rho\left(\phi\left(\bm{X};{\bm{S}},\mathcal{H}\right)\phi\left(\bm{X};{\bm{S}},\mathcal{H}\right)^{T}\right),\bm{S}\right),$ (6)
where $l\left(\cdot\right)$ is the binary cross entropy (BCE) and $\rho\left(\cdot\right)$ is the logistic function.
4.2 Learning Network Geometry with Generalized Graph Auto-encoders
The goal of the proposed framework is to learn a function that maps graphs to node representations and effectively match nodes from different graphs. This function is modeled by a GNN encoder $\phi\left(\bm{X};{\bm{S}},\mathcal{H}\right)$, where each layer is described by Eq. 4. The learned encoder should work for a family of training graphs $\left\{\mathcal{G}_{0},\dots,\mathcal{G}_{i},\dots,\mathcal{G}_{I}\right\}$ with a set of adjacency matrices $\mathbb{S}=\left\{\bm{S}_{0},\dots,\bm{S}_{i},\dots,\bm{S}_{I}\right\}$, rather than a single graph. So the idea is not to train an auto-encoder on a single graph but train a generalized graph auto-encoder by solving the following optimization problem.
@@ 4.2 Learning Network Geometry with Generalized Graph Auto-encoders
where $\bm{S}_{i}\in\mathbb{S}$ is a realization from a family of graphs and the expectation (empirical expectation is practice) is computed over this graph family. The generalized framework in (7) learns a mapping from graphs to node representations, and can be applied to out-of-distribution graphs that have not been observed during training. This twist in the architecture enables node embedding and graph matching for large-scale graphs, where training is computationally prohibitive.
@@ 4.3 Robust and Generalizable Node representations with self-supervised learning (data augmentation)
So far we proposed a convolutional framework to produce expressive node representations that are tailored to perform network alignment. In this subsection, we further upgrade our framework by ensuring the robustness and generalization ability of the proposed GNN mapping. In particular, for each graph, $\bm{S}_{i}\in\mathbb{S}$, we augment the training set with perturbed versions that are described by the following set of graph adjacencies $\mathbb{M}_{i}=\left\{\bm{S}_{i}^{(0)},\dots,\bm{S}_{i}^{(j)},\dots,\bm{S}_{i}^{(J)}\right\}$, that are perturbed versions of $\bm{S}_{i}$. To do so we add or remove an edge with a certain probability yielding $\tilde{\bm{S}}_{i}\in\mathbb{M}$, such that $\tilde{\bm{S}}_{i}=\bm{S}_{i}+\bm{M}_{i}$, where $\bm{M}_{i}\in\{-1,0,1\}^{N\times N}$. Note that $\bm{M}$ changes for each $\tilde{\bm{S}}_{i}$, and $\bm{M}[m,n]$ can be equal to $1$ and $-1$ only if $\bm{S}[m,n]$ is equal to 00 and $1$ respectively. To train the proposed generalized graph-autoencoder we consider the following optimization problem:
$\small\min_{\mathcal{H}}~{}\mathbb{E}_{\mathbb{S}}\left[\mathbb{E}_{\mathbb{M}_{i}}\left[l\left(~{}\rho\left(\phi\left(\bm{X};\tilde{\bm{S}}_{i},\mathcal{H}\right)\phi\left(\bm{X};\tilde{\bm{S}}_{i},\mathcal{H}\right)^{T}\right),\bm{S}_{i}\right)\right]\right],$ (8)
where $\mathbb{E}_{\mathbb{S}}$ is the expectation with respect to the family of graphs $\mathbb{S}$ and $\mathbb{E}_{\mathbb{M}_{i}}$ is the expectation with respect to the perturbed graphs $\mathbb{M}_{i}$. In practice, $\mathbb{E}_{\mathbb{S}},~{}\mathbb{E}_{\mathbb{M}}$ correspond to empirical expectations. Note that training according to (8) also benefits the robustness of the model, which is crucial in deep learning tasks (Wang et al., 2022). A schematic illustration of the training process can be found in Fig. 1.
@@
(Large-scale network alignment by transference)
The proposed framework learns a mapping $\phi:\mathbb{G}\to\mathbb{R}^{N\times F}$ that produces expressive and robust node representations for a family of graphs $\mathcal{G}\in\mathbb{G}$. This mapping is designed in such a way that the problem in (2) approximates the problem in (1) and allows solving network alignment in polynomial time. One of the main benefits of the proposed framework is that it enables large-scale network alignment. The transferability analysis of GNN encoders (Ruiz et al., 2020), suggests that we can train with small graphs and efficiently execute with much larger graphs when the substructures (motifs) that appear in the tested graphs, were also partially observed during training. Since the proposed generalized graph auto-encoder is trained with multiple graphs, a variety of motifs are observed during training, which cannot be observed with a classical graph autoencoder, and the proposed GNN encoder can be transferred to large-scale graphs.

### v4

# T-GAE: Transferable Graph Autoencoder for Network Alignment

@@ Abstract
Network alignment is the task of establishing one-to-one correspondences between the nodes of different graphs. Although finding a plethora of applications in high-impact domains, this task is known to be NP-hard in its general form. Existing optimization algorithms do not scale up as the size of the graphs increases. While being able to reduce the matching complexity, current GNN approaches fit a deep neural network on each graph and requires re-train on unseen samples, which is time and memory inefficient. To tackle both challenges we propose T-GAE, a transferable graph autoencoder framework that leverages transferability and stability of GNNs to achieve efficient network alignment on out-of-distribution graphs without retraining. We prove that GNN-generated embeddings can achieve more accurate alignment compared to classical spectral methods. Our experiments on real-world benchmarks demonstrate that T-GAE outperforms the state-of-the-art optimization method and the best GNN approach by up to 38.7% and 50.8%, respectively, while being able to reduce 90% of the training time when matching out-of-distribution large scale networks. We conduct ablation studies to highlight the effectiveness of the proposed encoder architecture and training objective in enhancing the expressiveness of GNNs to match perturbed graphs. T-GAE is also proved to be flexible to utilize matching algorithms of different complexities. Our code is available at https://github.com/Jason-Tree/T-GAE.
@@ 1 Introduction
Network alignment, also known as graph matching, is a classical problem in graph theory, that aims to find node correspondence across different graphs and is vital in a number of high-impact domains (Emmert-Streib et al., 2016). In social networks, for instance, network alignment has been used for user deanonymization (Nilizadeh et al., 2014) and analysis (Ogaard et al., 2013), while in bioinformatics it is a key tool to identify functionalities in protein complexes (Singh et al., 2008), or to identify gene–drug modules (Chen et al., 2018a). Graph matching also finds application in computer vision (Conte et al., 2003), sociology (Racz and Sridhar, 2021), to name a few. However, this problem is usually cast as a quadratic assignment problem (QAP), which is in general NP-hard.
@@
Various approaches have been developed to tackle network alignment and can be divided into two main categories; i) optimization algorithms that attempt to approximate the QAP problem by relaxing the combinatorial constraints, ii) embedding methods that approach the problem by implicitly or explicitly generating powerful node embeddings that facilitate the alignment task. Optimization approaches, as (Anstreicher and Brixius, 2001; Vogelstein et al., 2015) employ quadratic programming relaxations, while (Klau, 2009) and (Peng et al., 2010) utilize semidefinite or Lagrangian-based relaxations respectively, (Du et al., 2019) and (Du et al., 2022) proposed to solve network alignment together with link prediction. Successive convex approximations were also proposed by (Konar and Sidiropoulos, 2020) to handle the QAP. Challenges associated with these methods include high computational cost, infeasible solutions, or nearly optimal initialization requirements. Embedding methods, on the other hand, overcome these challenges, but they usually produce inferior solutions, due to an inherent trade-off between embedding permutation-equivariance and the ability to capture the structural information of the graph. Typical embedding techniques include spectral and factorization methods (Umeyama, 1988; Feizi et al., 2019; Zhang and Tong, 2016; Kanatsoulis and Sidiropoulos, 2022), structural feature engineering methods (Berlingerio et al., 2013; Heimann et al., 2018), and random walk approaches (Perozzi et al., 2014; Grover and Leskovec, 2016a). Recently (Chen et al., 2020; Karakasis et al., 2021) have proposed joint node embedding and network alignment, to overcome these challenges, but these methods do not scale up as the size of the graph increases.
Figure 1: To enhance the expressiveness of GNNs to align the unseen graphs, our proposed encoder processes the input features by a local MLP, we then incorporate attention mechanism on (1) input to each message-passing layer by attending to the output of the last layer and processed input feature. (2) output of the encoder by attending to the output of each message passing layer.
Graph Neural Networks (GNNs) are powerful architectures that learn graph representations (embeddings) in a self-supervised way (Kipf and Welling, 2016; You et al., 2020). They have shown state-of-the-art performance in several tasks, including biomedical (Gainza et al., 2020; Strokach et al., 2020; Jiang et al., 2021; Hu et al., 2023; Li et al., 2024), quantum chemistry (Gilmer et al., 2017), social networks and recommender systems (Ying et al., 2018; Wu et al., 2020; Liu et al., 2023a; Yang et al., 2024). A line of studies has been conducted to formulate key-point matching on images as graph matching problems (Fey et al., 2020; Yu et al., 2020). These frameworks rely on a powerful domain-specific encoder (CNNs, for example) to provide high-quality features. Given these high-quality features, GNNs are able to match graphs without training (Liu et al., 2022). However, expressive features are expensive and sometimes infeasible to build (Zhang et al., 2020; Zhou et al., 2021). (Liang et al., 2021) uses another trainable matrix to parameterize the internal connectivity between nodes of different graphs for faithful node alignment. Recently, (Gao et al., 2021) proposed use GNNs to learn node embedding, and match nodes with small Wasserstein distances. However, this method needs to fit a GNN on every input graph, which results in very high training cost, since training deep GNNs with large sizes graphs is computationally prohibitive, as GNNs have limited scalability with respect to graph and model sizes (Chen et al., 2018b, c; Chiang et al., 2019; Zeng et al., 2020).
To address these challenges, we propose T-GAE, a novel self-supervised GNN framework to perform network alignment. Specifically, we propose to utilize the transferability and robuseness of GNN to produce permutation equivariant and highly expressive embeddings. T-GAE trains the encoder on multiple families of small graphs and produce expressive/permutation equivariant representations for larger unseen networks. We further prove that GNN representations combine the eigenvectors of the graph in a nonlinear fashion and are at least as good in network alignment as certain spectral methods. T-GAE is a one-shot solution that tackles the challenges of real-time network alignment from (1) Optimization based algorithms: high computational cost, assume ground truth node correspondence as initialization. (2) Deep-learning based frameworks: re-train for every pair of graphs, rely on high quality of node features. Extensive experiments with real-world benchmarks demonstrate the effectiveness and efficiency of the proposed approach in both graph and sub-graph matching, thereby sheds light on the potential of GNN to tackle the highly-complex network optimization problems.
@@ 2 Preliminaries
Graphs are represented by $\mathcal{G}:=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{1,\dots,N\}$ is the set of vertices (nodes) and =$\mathcal{E}=\left\{\left(v,u\right)\right\}$ correspond to edges between pairs of vertices. A graph is represented in a matrix form by a graph operator $\bm{S}\in\mathbb{R}^{N\times N}$, where $\bm{S}(i,j)$ quantifies the relation between node $i$ and node $j$ and $N=|\mathcal{V}|$ is the total number of vertices. In this work, we use the normalized graph adjacency and study the most general form of network alignment where there is no given graph attributes and ground truth node correspondence(anchor links).
@@ 2.1 Network Alignment
Figure 2: Proposed training objective: For each of the training graph, we generate a number of augmented samples by randomly adding or removing edges. The node embedding of these augmented samples are decoded non-parametrically, and compared with the corresponding original graph to train the T-GAE encoder.
@@ 2.2 Spectral Decomposition of the Graph
where $\bm{E}\in\mathbb{R}^{N\times F}$ is embedding matrix and $\bm{E}[i,:]$ is the vector representation of node $i$. The optimization problem in (2) is a linear assignment problem and can be optimally solved in $\mathcal{O}\left(N^{3}\right)$ by the Hungarian method (Kuhn, 1955a). Simpler sub-optimal alternatives also exist that operate with $\mathcal{O}\left(N^{2}\right)$ or $\mathcal{O}\left(N~{}\text{log}(N)\right)$ flops.
@@ 3 Graph Neural Networks (GNNs) Upper-Bounds Spectral Methods for Network Alignment
The above property is not satisfied by other spectral methods. GNNs are also stable (Gama et al., 2020; Parada-Mayorga et al., 2023; Ruiz et al., 2021), transferable (Ruiz et al., 2020), and have high expressive power (Xu et al., 2019a; Abboud et al., 2021; Kanatsoulis and Ribeiro, 2022).
@@ 3.1 GNNs and Network Alignment
To characterize the ability of a GNN to perform network alignment we first pointed out the GNNs perform nonlinear spectral operations. Details can be found in Appendix Section D.1. We can further prove that:
@@
The proof can be found in Appendix D. The assumption that the graph adjacencies have different eigenvalues is not restrictive. Real nonisomorphic graphs have different eigenvalues with very high probability (Haemers and Spence, 2004). Theorem 3.2 compares the network alignment power of a GNN with that of a spectral algorithm Umeyama (1988), that uses the absolute values of graph adjacency eigenvectors to match two different graphs. According to Theorem 3.2 there always exists a GNN that can perform at least as well as the spectral approach. The proof studies a GNN with white random input and measures the variance of the filter output. Then it shows that message-passing layers are able to compute the absolute values of the graph adjacency eigenvectors when the adjacency has non-repeated eigenvalues. As a result there always exists a single layer GNN that outputs the same node features as the ones used in Umeyama (1988), which concludes our proof. The questions is: How do we train such a GNN that is (1) Expressive to the structural information thus its output can be used to match corresponding nodes. (2) Robust to different perturbations and even larger scale unseen graphs so that it can be deployed efficiently without re-training. (3) Agnostic to the ground truth node correspondence so that it can be trained unsupervisedly, which makes it generalizable to most real-world settings. To answer these questions, we introduce our proposed training framework in the following section.
@@ 4 Proposed Method
We now leverage the favorable properties of GNNs (permutation equivariance, expressivity, and transferability) to tackle real-world network alignment. Our approach learns low-dimensional node embeddings (Eq. 4) that enable graph matching via solving the linear assignment in (2) rather than a quadratic assignment problem in (1). We design a robust GNN framework such that the node embeddings are expressive to accurately match similar nodes and also stable to graph perturbations.
4.1 Learning Network Geometry with Transferable Graph Auto-encoders
The goal of the proposed framework is to learn a function that maps graphs to node representations and effectively match nodes from different graphs. This function is modeled by a GNN encoder $\phi\left(\bm{X};{\bm{S}},\mathcal{H}\right)$, described by Fig. 1. The learned encoder should work for a family of training graphs $\left\{\mathcal{G}_{0},\dots,\mathcal{G}_{i},\dots,\mathcal{G}_{I}\right\}$ with a set of adjacency matrices $\mathbb{S}=\left\{\bm{S}_{0},\dots,\bm{S}_{i},\dots,\bm{S}_{I}\right\}$, rather than a single graph. So the idea is not to train a GNN on a single graph (Kipf and Welling, 2016), but train a transferable graph auto-encoder by solving the following optimization problem.
@@ 4.1 Learning Network Geometry with Transferable Graph Auto-encoders
where $l\left(\cdot\right)$ is the binary cross entropy (BCE) and $\rho\left(\cdot\right)$ is the logistic function. $\bm{S}_{i}\in\mathbb{S}$ is a realization from a family of graphs and the expectation (empirical expectation is practice) is computed over this graph family. The generalized framework in (6) learns a mapping from graphs to node representations, and can be applied to out-of-distribution graphs that have not been observed during training. This twist in the architecture enables node embedding and graph matching for the unseen and larger scale networks without re-training, where fitting a GNN is computationally prohibitive in real-world applications.
@@ 4.2 Robust and Generalizable Node representations with self-supervised learning (data augmentation)
So far we proposed a GNN framework to produce expressive node representations to perform network alignment. In this subsection, we further upgrade our framework by ensuring the robustness and generalization ability of the proposed mapping. In particular, for each graph, $\bm{S}_{i}\in\mathbb{S}$, we augment the training set with perturbed versions that are described by the following set of graph adjacencies $\mathbb{M}_{i}=\left\{\bm{S}_{i}^{(0)},\dots,\bm{S}_{i}^{(j)},\dots,\bm{S}_{i}^{(J)}\right\}$, that are perturbed versions of $\bm{S}_{i}$. To do so we add or remove an edge with a certain probability yielding $\tilde{\bm{S}}_{i}\in\mathbb{M}$, such that $\tilde{\bm{S}}_{i}=\bm{S}_{i}+\bm{M}_{i}$, where $\bm{M}_{i}\in\{-1,0,1\}^{N\times N}$. Note that $\bm{M}$ changes for each $\tilde{\bm{S}}_{i}$, and $\bm{M}[m,n]$ can be equal to $1$ and $-1$ only if $\bm{S}[m,n]$ is equal to 00 and $1$ respectively. To train the proposed transferable graph-autoencoder we consider the following optimization problem:
$\min_{\mathcal{H}}~{}\mathbb{E}_{\mathbb{S}}\left[\mathbb{E}_{\mathbb{M}_{i}}\left[l\left(~{}\rho\left(\phi\left(\bm{X};\tilde{\bm{S}}_{i},\mathcal{H}\right)\phi\left(\bm{X};\tilde{\bm{S}}_{i},\mathcal{H}\right)^{T}\right),\bm{S}_{i}\right)\right]\right],$ (7)
where $\mathbb{E}_{\mathbb{S}}$ is the expectation with respect to the family of graphs $\mathbb{S}$ and $\mathbb{E}_{\mathbb{M}_{i}}$ is the expectation with respect to the perturbed graphs $\mathbb{M}_{i}$. In practice, $\mathbb{E}_{\mathbb{S}},~{}\mathbb{E}_{\mathbb{M}}$ correspond to empirical expectations. Note that training according to (7) also benefits the robustness of the model, which is crucial in deep learning tasks (Wang et al., 2022; He, 2021). A schematic illustration of the training process can be found in Fig. 2.
@@
(Large-scale network alignment by transferability of GNN)
The proposed framework learns a mapping $\phi:\mathbb{G}\to\mathbb{R}^{N\times F}$ that produces expressive and robust node representations for a family of graphs $\mathcal{G}\in\mathbb{G}$. This mapping is designed in such a way that the problem in (2) approximates the problem in (1) and allows solving network alignment in polynomial time. One of the main benefits of the proposed framework is that it enables large-scale network alignment. Task specific augmentation during training is the key to prompt transferability of deep neural networks (Li et al., 2022). And the transferability analysis of GNN encoders (Ruiz et al., 2020) suggests that we can train with small graphs and efficiently execute with much larger graphs when the substructures (motifs) that appear in the tested graphs, were also partially observed during training. Since the proposed transferable graph auto-encoder is trained with multiple graphs, a variety of motifs are observed during training, which cannot be observed with a classical graph autoencoder, and the proposed GNN encoder can be transferred to larger-scale out-of-distribution graphs.